Fermi National Laboratory


Accelerator Update for March 9, 2005 - March 11, 2005

Wednesday March 9
The day shift began with the Tevatron (TeV) in shot setup, with Switchyard (SY) and NuMI beam off due to shot setup, and with MiniBooNE taking beam.

At 8:40 AM, Operations reported that the shot setup was on hold due to TeV experts adjusting the tune, chromaticity, orbit, and the helix.

At 9:31 AM, right after the antiprotons were loaded into the TeV, the TeV suffered a multi-house quench in sectors A, B, and F.

The Antiproton Source (Pbar) resumed stacking at 9:45 AM.

Operators readied the Collider Detector at Fermilab's (CDF) collision hall for access at 10:19 AM.

Operations established beam to the Meson experiments MCenter and MTest at 1:05 PM.

A Main Injector (MI) expert began a bunch rotation study at 1:52 PM.

At 2:37 PM, another MI expert began a different study that halted SY beam.

Operations resumed sending beam to SY and the Meson experiments around 4:50 PM.

Operations put the TeV into a dry squeeze at 6:18 PM.

Operations loaded a 36x0 study-store into the TeV at 8 PM and then prepared the Recycler to stash antiprotons.

At 8:22 PM, the TeV study-store aborted due to corrupted data.

At 8:54 PM, Operations tried to resume stacking when all the TeV kickers fired. Experts began their investigation of the event.

Pbar resumed stacking at 8:57 PM.

At 9:38 PM, the TeV beam aborted due to the power supply for the defocusing sextupoles tripping off. An expert was called in.

Operations put the TeV into a dry squeeze at 9:48 PM.

MCenter and MTest resumed taking beam at 9:51 PM.

NuMI experts said they were finished using the beam for the night at 11:49 PM.

Thursday March 10
The midnight shift began with the TeV conducting studies, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

At 12:24 AM, Operations reported that the sector F17 collimator wasn't moving to its safe position. Operations contacted an expert.

At 1:14 AM, the Booster East Anode power supply tripped. All attempts to reset it failed. Operators called in experts.

Operations put the TeV through a dry squeeze at 1:24 AM.

Booster experts discovered at 2:26 PM that a RF station (BRF1) was keeping the Anode power supply off. They bypassed the station. Booster resumed accelerating beam and Pbar resumed stacking.

Operations loaded the TeV with a 36x0 study-store at 3:25 AM.

At 3:39 AM, the TeV suffered a sector F17 quench due to a collimator problem. Operators called in an expert to help investigate the problem.

MTest experimenters reported at 4:01 AM that they were finished with the beam for the night.

At 7:47 AM, two MI RF stations (MIRF2 & 7) tripped off. MIRF7 wouldn't come on. Operators contacted an expert.

The Neutron Therapy Facility (NTF) requested beam for patient treatment at 8:01 AM.

Operations reported at 8:38 AM that the F17 collimator work was finished. Operations prepared the TeV for a 12x0 study-store.

NuMI experts began a study period at 8:39 AM.

At 9:30 AM, the TeV suffered a quench at sector F1. Experts investigated.

Operations began a TeV checkout with only a QBS ramp at 11:32 AM.

Booster experts began their work on GMPS at 12:17 PM.

Booster RF experts repaired BRF1 at 1:15 PM.

GMPS work finished at 2:32 PM.

MiniBooNE resumed taking beam at 2:43 PM.

Pbar resumed stacking at 2:49 PM.

NuMI experts resumed their study period at 3:02 PM. MTest also resumed taking beam.

At 3:13 PM, BRF17 & 18 tripped off. An operator reset the stations, but then BRF17 tripped on a ground fault. An expert was asked to investigate.

At 4:08 PM, BRF17 tripped off. It reset okay.

At 4:18 PM, most of the MI RF stations tripped off. An expert saw this and reset them. He didn't know why they had tripped.

Operations reported at 5:47 PM that they had loaded a 12x0 study-store into the TeV.

Operations put the TeV into shot setup at 6:16 PM.

Operations established store 4038 at 8:55 PM with an initial luminosity of 95.9E30. Pbar resumed stacking.

NuMI studies resumed at 9:02 PM.

At 11:32 PM, the TeV quenched at sectors C4 and D0. Operators investigated.

NuMI studiers finished for the night at 11:46 PM.

Friday March 11
The midnight shift began with operators investigating the TeV quench, and with MiniBooNE , MCenter, and MTest taking beam.

Operators reported at 12:40 AM that they found some D0 low beta voltage spikes on the Quench Protection Monitor (QPM) system and reported them to experts who continued the investigation.

A controls expert joined another expert at 2:36 AM to help with the quench investigation.

At 5:02 AM, the Linac suffered a trip.

A controls expert began replacing the D0 QPM at 3:40 AM. He finished the job at 5:09 AM.

At 5:12 AM, Operations reported that they haven't been able to restore beam to Booster since the Linac trip. An operator called in an expert.

Operations put the TeV through a dry squeeze at 5:50 PM.

Operations reported at 6:18 AM that the controls experts were going to do more QPM work.

A controls expert reported at 6:47 AM that the D0 QPM noise hadn't gone away. Investigations continued.

Operations reported at 7:39 AM that the Booster work was completed and that beam had been established to the MI dump. Pbar resumed stacking at 7:44 AM.

The Plan for Monday
The plan for today is to resolve the QPM noise problem and then put the TeV into a long dry squeeze before putting the TeV into shot setup.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/11/2005   email Fermilab